# Learning an Image Editing Model without Image Editing Pairs

### v1

@@ Appendix B Ablation Study
Figure 9: Limitation. Our method can struggle to maintain exact pixel consistency between the input and edited image. Having LPIPS (Zhang et al., 2018) loss between the input and output edited image can resolve it to an extent (top row) but at the cost of reduced editing success (bottom row).
@@ Appendix C Limitation
A limitation of our framework is the lack of pixel-wise supervision to preserve regions that should remain unchanged under a given edit instruction. Consequently, edited images can deviate from the input image in details or spatial alignment, as shown in Figure 9, $1^{\text{st}}$ column. While our VLM-based editing loss includes a question to check consistency between the input and edited images, it does not enforce pixel-level alignment. Empirically, we find that current VLMs struggle to detect subtle changes. To mitigate this, we experiment with an additional LPIPS (Zhang et al., 2018) loss between the input and output edited images. As shown in the last column of Figure 9, this improves consistency but also negatively impacts editing quality, particularly for edit-types like Removal. Future work could explore specialized VLMs that are more sensitive to fine-grained, pixel-level differences.
@@ E.1 Local-image editing
Training hyperparameters. We train on a batch-size of $32$ using Adam (Adam et al., 2014) optimizer with a learning rate of $2\times 10^{-6}$, $\beta_{1}$ as 0, and $\beta_{2}$ as $0.9$. We train for a total of 1010K iteratuions with auxiliary network, $A_{\phi}$, being updated $10$ times for every generator, $G_{\theta}$, update, following similar strategy adopted in DMD2 (Yin et al., 2024a). We train with the identity loss (Section 4.3) for $250$ iterations. For faster convergence, the first 44K training iterations are trained with a single step prediction ($t=1$ in Line 3 of Algorithm 1), and then we start the 2-step unrolling of the diffusion trajectory. The final loss is a weighted combination of VLM-based editing loss and distribution matching loss with $\lambda_{\text{VLM}}=0.01$ and $\lambda_{\text{DMD}}=0.5$. During training, we also add a “do nothing” editing task with $\mathcal{L}_{2}$ loss between the input and edited image as regularization with a $1\%$ probability. This helps the model learn to maintain consistency between input and edited images. During training, we sample the editing instruction corresponding to each subtype uniformly, except removal, which is sampled with $25\%$ probability. This is because, empirically, we observe that removal is more difficult than other edit-types like Color change.
@@ Appendix F Other Baseline Details
Flow-GRPO (Liu et al., 2025a). We follow the open-source implementation of Flow-GRPO and train with the same computational budget as our method, i.e., across $4$ A100 GPUs and $2.5$ days of training. The final model is fine-tuned from a pre-trained image-editing model for 55K iterations. During training, we collect $16$ images per-prompt with $12$ denoising steps ($28$ during inference) for computing the mean and standard deviation in GRPO (Shao et al., 2024). Following their official implementation, we train with LoRA (Hu et al., 2022) of rank $32$, $\alpha=64$, learning rate $1\times 10^{-4}$, and use the VLM to score edits on a scale of 0 to 9, which is normalized between 0-1 to get the final reward. The exact prompt used to query the VLM is derived from VIEScore (Ku et al., 2024) and is shown below.
@@
Sampling parameters for customization baselines. Here as well, we follow the open-source implementation to sample images from all the baseline models for the benchmark evaluations. In the case of DSD (Cai et al., 2025), it employs Gemini-1.5 to convert the input user-prompt into a detailed prompt. However, we skipped this step for a fair evaluation with other methods, which do not use any prompt rewriting tools. In the case of SynCD (Kumari et al., 2025), though it supports multiple reference images as input, we evaluated it with a single reference image, to keep the setup similar to other baseline methods and Ours. For sampling images with OminiControl (Tan et al., 2025) and DSD (Cai et al., 2025), we follow their recommended prompt setting and replace the category name with “this item”.
Appendix G Societal Impact

### v2

@@ Appendix B Ablation Study
[…]
Figure 10: Performance for each sub edit-type for LPIPS and cycle-loss on GEdit-Bench. Additional regularization losses like LPIPS between input and edited image or cycle loss help increase the alignment between input and output edited image as shown in Figure 11, but lower the overall editing success.
@@ Appendix C Limitation
Alignment between input and edited image. A limitation of our framework is the lack of pixel-wise supervision to preserve regions that should remain unchanged given an edit instruction. Consequently, edited images can deviate from the input image in details or spatial alignment, as shown in Figure 11, $1^{\text{st}}$ column. While our VLM-based editing loss includes a question to check consistency between the input and edited images (Identity-preservation question in Section 4.2), it does not enforce pixel-level alignment, since current VLMs struggle to detect subtle changes and minor details.
Here, we experiment with two regularization losses that can improve the alignment between input and edited images in regions that should not be edited: (1) LPIPS (Zhang et al., 2018) loss between input and edited image, and (2) Cycle loss (Zhu et al., 2017) between input and cycle-edited image.
In the cycle loss, for an edited image (e.g., remove the flower pot), we apply a reverse edit to it (e.g., add a flower pot) and encourage this reversal to reconstruct the original input image. We compute the reconstruction loss between the input and reverse-edited image in the SigLIP (Zhai et al., 2023) feature space. We select SigLIP’s global embedding over fine-grained feature spaces like LPIPS to provide flexibility when reverse instructions are ambiguous (e.g., restoring the exact original object). For training with cycle loss, we add it to our training objective after 44K iterations. The reverse edit instruction for each (image, instruction) pair is generated using Qwen-32B VLM.
While LPIPS loss improves consistency, it degrades editing quality, particularly for edit-types like Removal, as shown in Figure 11. Cycle loss works better in comparison but still has lower edit success than ours, shown in Figure 10. The overall VIEScore on GEdit-Bench for LPIPS and cycle loss is $4.52$ and $5.21$ respectively, compared to $6.10$ for Ours. However, we anticipate that as VLMs improve, with the capability of nuanced, pixel-level assessment, it will also help improve the input-edit alignment in our method without requiring additional regularization.
Figure 11: Limitation. We use an Identity-preservation question, Section 4.2, to maintain the consistency between input and edited image ($1^{\text{st}}$ vs. $2^{\text{nd}}$ column). However, given the lack of pixel-level supervision, it can still struggle in many scenarios to maintain exact identity. Adding cycle loss between input and cycle-edited image helps mitigate this issue to an extent ($3^{\text{rd}}$ column), though with a drop in performance (Overall VIEscore of $5.10$ vs $6.10$ for Ours). We also experimented with LPIPS (Zhang et al., 2018) loss between the input and edited image, which helps ($1^{\text{st}}$ row) but at the cost of completely failing on tasks like removal ($2^{\text{nd}}$ row).
VLM supervision for difficult editing tasks. While NP-Edit can successfully edit images for various challenging edit-types like actions ($2^{\text{nd}}$ last row in Figure 2) and shape changes (Figure 12 (a)), our method is fundamentally upper bounded by the VLM’s ability to understand and reason about image editing. Consequently, it currently fails at tasks requiring complex spatial reasoning (e.g., moving objects), as VLMs struggle to provide meaningful feedback in these scenarios. We show sample failure case examples in Figure 12 (b). As unified generative models and Vision–Language Models advance, the instruction-following capabilities learned through our NP-Edit training framework are expected to improve correspondingly.
Figure 12: Challenging and difficult editing tasks. Our method can perform various challenging edits as well, like shape change as shown in (a). But we fail at other difficult tasks requiring spatial reasoning, such as moving objects, where VLM-based gradient feedback is less reliable.
@@ E.1 Local-image editing
Training hyperparameters. We train on a batch-size of $32$ using Adam (Adam et al., 2014) optimizer with a learning rate of $2\times 10^{-6}$, $\beta_{1}$ as 0, and $\beta_{2}$ as $0.9$, on $32$ A100-80GB GPUs. The few-step sampling time schedule, $t_{4},t_{3},t_{2},t_{1}]$ is set to $[1.0,0.90,0.70,0.47]$, shifted towards more noisy region (Esser et al., 2024). We train for a total of 1010K iterations with $N_{\text{aux}}=10$, i.e., auxiliary network, $A_{\phi}$, being updated $10$ times for every generator, $G_{\theta}$, update, following similar strategy adopted in DMD2 (Yin et al., 2024a). We train with the identity loss (Section 4.3) for $250$ iterations. For faster convergence, the first 44K training iterations are trained with a single step prediction ($t=1$ in Line 3 of Algorithm 1), and then we start the 2-step unrolling of the diffusion trajectory. The final loss is a weighted combination of VLM-based editing loss and distribution matching loss with $\lambda_{\text{VLM}}=0.01$ and $\lambda_{\text{DMD}}=0.5$. During training, we also add a “do nothing” editing task with $\mathcal{L}_{2}$ loss between the input and edited image as regularization with a $1\%$ probability. This helps the model learn to maintain consistency between input and edited images. During training, we sample the editing instruction corresponding to each subtype uniformly, except removal, which is sampled with $25\%$ probability. This is because, empirically, we observe that removal is more difficult than other edit-types like Color change.
@@ Appendix F Other Baseline Details
Flow-GRPO (Liu et al., 2025a). We follow the open-source implementation of Flow-GRPO and train with the same computational budget as our method, i.e., across $4$ A100 GPU nodes (32 GPUs) and $2.5$ days of training. The final model is fine-tuned from a pre-trained image-editing model for 55K iterations. During training, we collect $16$ images per-prompt with $12$ denoising steps ($28$ during inference) for computing the mean and standard deviation in GRPO (Shao et al., 2024). Following their official implementation, we train with LoRA (Hu et al., 2022) of rank $32$, $\alpha=64$, learning rate $1\times 10^{-4}$, and use the VLM to score edits on a scale of 0 to 9, which is normalized between 0-1 to get the final reward. The exact prompt used to query the VLM is derived from VIEScore (Ku et al., 2024) and is shown below.
@@
Sampling parameters for customization baselines. We follow the open-source implementation to sample images from all the baseline models for the benchmark evaluations. In the case of DSD (Cai et al., 2025), it employs Gemini-1.5 to convert the input user-prompt into a detailed prompt. However, we skipped this step for a fair evaluation with other methods, which do not use any prompt rewriting tools. In the case of SynCD (Kumari et al., 2025), though it supports multiple reference images as input, we evaluated it with a single reference image, to keep the setup similar to other baseline methods and Ours. For sampling images with OminiControl (Tan et al., 2025) and DSD (Cai et al., 2025), we follow their recommended prompt setting and replace the category name with “this item”.
Appendix G Use of LLM in Paper Writing.
We primarily use ChatGPT as a grammatical aid and for polishing paragraphs, instructing it to “refine the text without adding or omitting any relevant information.”
Appendix H Reproducility of Results
We provide full implementation details of our method, including dataset construction, training algorithm, and hyperparameters in Section 4 and Appendix D, E. A pseudo-code description is also included in Algorithm 1. Together, these are intended to facilitate the reproducibility of all results reported in our paper.
Appendix I Ethics Statement
@@
Figure 16: Qualitative comparison on TED-Bench. We show results of our and baseline image-editing methods under the few-step sampling setting. For comparison, we also show the results of the best method with multi-step sampling, as measured by the quantitative metrics (Table 8), in the $1^{\text{st}}$ column. Our method performs on par or better than baseline methods across different edit types in the few-step setting.